OpenStudy (anonymous):

I need help understanding this question:Determine if the event shows uniform probability or not. You roll two die, one a standard die, and the other a die with just the even numbers between 1 and 6. I will give a medal if you can help me! :)

OpenStudy (asevilla5):

if by "uniform probability", you mean that each event has the same probability, 1) YES, Pr = 1/6*1/3 = 1/18 for each event